A force of magnitude 200N is directed west, and a second force of magnitude 100N is directed east of north. Find the magnitude and direction of a third force that, when added to the first two forces, gives a resultant whkse magnitude is zero


1
Expert's answer
2020-12-10T11:10:23-0500

Using formula R = (R1+ R22 - 2*R1*R2*cos"\theta" )0.5, net magnitude of can be obtained.

"\theta" = (90+45)"\degree" in this case.

Fr is resultant force initially applied.

Fo is opposing force needed to make net force 0.

The Resultant Direction of first two forces, Fr, would be west of north.

So, The direction of third force, Fo, would be East of South.

Magnitude of Fwould be 147.36 N